# Positive definiteness and Fell bundles over discrete groups

### v1

@@ Abstract
We introduce a natural concept of positive definiteness for bundle maps between Fell bundles over (possibly different) discrete groups and describe several examples. Such maps induce completely positive maps between the associated full cross-sectional C*-algebras in a functorial way. Under the assumption that the kernel of the homomorphism connecting the groups under consideration is amenable, they also induce completely positive maps between the associated reduced cross-sectional C*-algebras. As an application, we define an approximation property for a Fell bundle over a discrete group which generalizes Exel’s approximation property and still implies the weak containment property; both approximation properties coincide when the unit fibre is nuclear.
@@ 1 Introduction
The concept of positive definiteness for complex-valued functions on a group $G$ has a long history (it goes back to Toeplitz in 1911 when $G={\mathbb{Z}}$ and to Mathias in 1923 when $G={\mathbb{R}}$; for more details, see [59] and references therein). Its importance is ubiquitous in various areas of mathematics, in particular in connection with the development in the 1940’s and 50’s of the representation theory of locally compact groups and of the associated group $C^{*}$-algebras (see e.g. [22, 28, 29]). On the other hand, completely positive linear maps play a prominent rôle in the modern theory of operator algebras and in quantum information theory (cf. e.g. [56, 14, 33]). Haagerup discovered in 1978 that there is a very useful link between the two concepts. Namely, he proved in [31, Theorem 3.1] that given an action $\alpha$ of a locally compact group $G$ on a von Neumann algebra ${\cal M}$, every continuous positive definite function $\psi$ on $G$ gives rise to a normal completely positive linear map $M_{\psi}$ on the crossed product ${\cal M}\rtimes_{\alpha}G$ such that $M_{\psi}$ acts on each $g$-component by multiplication with $\psi(g)$. This link also holds in the context of group $C^{*}$-algebras and group von Neumann algebras, as noted by Haagerup in his seminal paper [32] and exploited in later works, such as [21] where continuous positive definite functions on $G$ are identified as multipliers of the Fourier algebra $A(G)$ of $G$. In the setting of discrete unital $C^{*}$-dynamical systems, a similar link was established in a series of paper [8, 9, 10], where broader concepts of multipliers were introduced. As a sample of related articles we mention [50, 51, 52, 34, 6, 39, 49].
@@
Given two $C^{*}$-algebras $A$ and $B$, it is often the case that one needs to construct completely positive linear maps from $A$ to $B$. When $A$ and $B$ belong to some class of $C^{*}$-algebras, one may hope to find some specific procedure achieving this. In this paper, we consider this problem for the class $\mathcal{F}$ (resp. $\mathcal{F}_{\rm r}$) of $C^{*}$-algebras which may be described as the full (resp. reduced) cross-sectional $C^{*}$-algebra $C^{*}({\mathcal{A}})$ (resp. $C_{r}^{*}({\mathcal{A}})$) of a Fell bundle ${\mathcal{A}}$ over a discrete group, as studied in [26, 27] (and e.g. in [45, 40]; see also [28, 29] for a more general treatment dealing with $C^{*}$-algebraic bundles, i.e., Fell bundles over locally compact groups). The class $\mathcal{F}$ (resp. $\mathcal{F}_{\rm r}$) is large as it contains all full (resp. reduced) twisted group $C^{*}$-algebras, and more generally all full (resp. reduced) $C^{*}$-crossed products associated to discrete twisted $C^{*}$-dynamical systems, even those for which the action of the group is only assumed to be partial [25]. Also, every $C^{*}$-algebra on which there exists a maximal (resp. normal) nondegenerate coaction by a discrete group belongs to $\mathcal{F}$ (resp. $\mathcal{F}_{\rm r}$), see [23, Proposition 4.2] (resp. [58, Corollary 3.9]). In particular, both classes include any $C^{*}$-algebra on which there exists a nondegenerate coaction by a discrete amenable group, e.g., any $C^{*}$-algebra on which there exists an action of a compact abelian group, such as Cuntz algebras and, more generally, (higher-rank) graph $C^{*}$-algebras. Note here that one may always choose to regard a $C^{*}$-algebra $B$ as the cross-sectional $C^{*}$-algebra of a Fell bundle over the trivial group, hence that our set up includes the case where no assumption is put on $B$.
@@
After a preliminary section (Section 2), our main findings are gathered in Section 3. To describe these, let us first assume that $A=C^{*}({\mathcal{A}})$ and $B=C^{*}({\mathcal{B}})$, where ${\mathcal{A}}$ is a Fell bundle over a discrete group $G$ and ${\mathcal{B}}$ is a Fell bundle over a discrete group $H$, and let $\varphi:G\to H$ be a homomorphism. Consider an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T:{\mathcal{A}}\to{\mathcal{B}}$, meaning that $T$ maps the fibre of ${\mathcal{A}}$ over any $g\in G$ into the fibre of ${\mathcal{B}}$ over $\varphi(g)$ in a linear and bounded way. Inspired by the definition of positive definiteness for a complex function on a group and the definition of complete positivity for a linear map between $C^{*}$-algebras, we introduce a natural notion of positive definiteness for such a bundle map, and show that $T$ is positive definite if and only if there exists a completely positive linear map $\Phi_{T}:A\to B$ which “extends” $T$ (see Theorem 3.10 for a more precise statement). Our proof relies on a reformulation of positive definiteness in terms of the matrix $C^{*}$ algebras associated to a Fell bundle by Abadie and Ferraro in [1], and builds on the Stinespring dilation theorem for Fell bundles due to Buss, Ferraro and Sehnem in [18, Appendix]. After introducing a category of Fell bundles over discrete groups whose arrows are pairs $(\varphi,T)$, where $\varphi\in{\rm Hom}(G,H)$ and $T$ is a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map, we get that the association ${\mathcal{A}}\mapsto C^{*}({\mathcal{A}}),\,(\varphi,T)\mapsto\Phi_{T}$ is a functor into the category of $C^{*}$-algebras with completely positive linear maps as arrows.
@@
Next, we consider the case where $A=C_{r}^{*}({\mathcal{A}})$ and $B=C_{r}^{*}({\mathcal{B}})$. It is not difficult to see that an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T$ has to be positive definite whenever there exists a completely positive linear map $M_{T}:A\to B$ which “extends” $T$. However, the converse statement can not hold in general. This may be deduced from a result in the recent book of Bekka and de la Harpe, cf. [11, Corollary 8.C.15], which says that the homomorphism $\varphi:G\to H$ extends to a *-homomorphism from $C_{r}^{*}(G)$ into $C_{r}^{*}(H)$ if and only if the kernel of $\varphi$ is an amenable subgroup of $G$. Using this result in combination with Fell’s absorption principle for Fell bundles and a generalization of another result of Buss, Ferraro and Sehnem ([18, Proposition 4.8]), we are able to show that the converse does hold if we also assume that $\ker(\varphi)$ is amenable, cf. Theorem 3.12. The association ${\mathcal{A}}\mapsto C_{r}^{*}({\mathcal{A}}),(\varphi,T)\mapsto M_{T}$ is then also functorial if one considers the subcategory of Fell bundles whose arrows satisfy that $\ker(\varphi)$ is amenable.
@@
In Section 4 we illustrate the usefulness of Theorem 3.10 and Theorem 3.12 in a long series of examples, showing that these two results are not only generalizations of previous results of the same flavor, but also lead to new applications. In particular, we devote a section (Section 5) to discuss a new approximation property for a Fell bundle ${\mathcal{A}}$ over a discrete group. We call it the PD-approximation property since it requires the existence of a net of uniformly bounded, finitely supported positive definite bundle maps converging to the identity map on each fibre, in analogy with the well-known characterization of the amenability of a discrete group in terms of positive definite functions. The PD-approximation property for ${\mathcal{A}}$ is (formally) weaker than Exel’s approximation property [26, 27], but still implies the weak containment property (i.e., amenability of ${\mathcal{A}}$ in the sense of Exel), cf. Theorem 5.4. Moreover, when it is satisfied, then $C^{*}({\mathcal{A}})$ (resp. $C^{*}_{r}({\mathcal{A}})$) is nuclear if and only if the unit fibre of ${\mathcal{A}}$ is nuclear, cf. Theorem 5.6. Using current knowledge from [3], it follows that the PD-approximation property for ${\mathcal{A}}$ is equivalent to Exel’s approximation property when the unit fibre of ${\mathcal{A}}$ is assumed to be nuclear. However, the arguments used to deduce this fact give no clue whether it also holds when the unit fibre of ${\mathcal{A}}$ is not nuclear.
@@
When we were about to finish the first draft of the present article, an interesting preprint by Buss, Kwasniewski, McKee and Skalski appeared [19], where the existing theories of Fourier-Stieltjes algebras associated to groups [36], to groupoids [60, 54] and to discrete twisted unital $C^{*}$-dynamical systems [10] is extended to the category of twisted actions by étale groupoids on $C^{*}$-bundles. Only a few of their results deal with the general case of Fell bundles over étale groupoids, hence with general Fell bundles over discrete groups, and there is little overlap between our papers.
@@ 2 Preliminaries
Let ${\mathcal{A}}=(A_{g})_{g\in G}$ be a Fell bundle over $G$. So each fibre $A_{g}$ is a Banach space and we often use ${\mathcal{A}}$ to also denote the disjoint union of the $A_{g}$’s. There is a multiplication map ${\mathcal{A}}\times{\mathcal{A}}\to{\mathcal{A}},(a,b)\mapsto ab$, and an involutive map ${\mathcal{A}}\to{\mathcal{A}},a\mapsto a^{*}$, such that $A_{g}A_{h}\subseteq A_{gh}$ and $A_{g}^{*}=A_{g^{-1}}$ for all $g,h\in G$, which satisfy some natural conditions and turn the unit fibre $A_{e}$ into a $C^{*}$-algebra. The vector space $C_{c}({\mathcal{A}})$ consisting of all finitely supported functions $f:G\to{\mathcal{A}}$ such that $f(g)\in A_{g}\text{ for all }g\in G$, becomes a *-algebra with respect to the operations defined by
@@
$\iota^{\mathcal{A}}(f)=\sum_{g\in G}\lambda^{\mathcal{A}}_{g}(f(g))\quad\text{ for all }f\in C_{c}({\mathcal{A}}).$
@@
As each $\widehat{j}^{\mathcal{A}}_{g}$ is isometric, this gives that
@@
$E_{g}^{\mathcal{A}}\big{(}\lambda^{\mathcal{A}}_{h}(a)\big{)}=\begin{cases}a,\quad\text{ if }g=h,\\ 0,\quad\text{ if}~g\neq h.\end{cases}$
For $x\in C_{r}^{*}({\mathcal{A}})$, $E_{g}^{\mathcal{A}}(x)$ may be thought of as the Fourier coefficient of $x$ at $g$. The map $E^{\mathcal{A}}:=E_{e}^{\mathcal{A}}$ is a faithful conditional expectation from $C_{r}^{*}({\mathcal{A}})$ onto $A_{e}$ (when one identifies $A_{e}$ with $\lambda_{e}^{\mathcal{A}}(A_{e})$).
We will need the following lemma several times. It follows readily from [27, Lemma 17.2]. For completeness, we give the proof.
@@
Proof.
Write $a=b^{*}b$ for some $b\in A_{e}$. Then $c^{*}ac=c^{*}b^{*}bc=(bc)^{*}bc\in(A_{e})^{+}$ and [27, Lemma 17.2] gives that
$c^{*}ac=~c^{*}b^{*}bc~\,\leq\,\|b\|^{2}\,c^{*}c=\|a\|\,c^{*}c.$
Hence $\|c^{*}ac\|\,\leq\,\|a\|\,\|c^{*}c\|=\|a\|\,\|c\|^{2}$. $\square$
@@
is a $C^{*}$-algebra with respect to the natural operations, the norm being inherited from the natural embedding of $M_{\mathbf{g}}({\mathcal{A}})$ into $M_{n}(C^{*}({\mathcal{A}}))$ (or $M_{n}(C_{r}^{*}({\mathcal{A}}))$). We note that $M_{\mathbf{g}}({\mathcal{A}})$ can also be represented by adjointable operators on a right Hilbert $A_{e}$-module. Indeed, consider the right Hilbert $A_{e}$-module $A_{\mathbf{g}}:=A_{g_{1}^{-1}}\oplus\cdots\oplus A_{g_{n}^{-1}}$ obtained by taking the direct sum of the $A_{g_{i}}$’s (considered as right Hilbert $A_{e}$-modules), whose inner product is given by
@@ 3.1 Positive definite bundle maps
Considering the case where $G=H$ and $\varphi={\rm id}_{G}$, Exel defines a morphism from ${\mathcal{A}}$ to ${\mathcal{B}}$ to be an ${\mathcal{A}}$$-{\mathcal{B}}$ bundle map $T$ which is self-adjoint and multiplicative, and he proves that it gives rise to a *-homomorphism from $C^{*}({\mathcal{A}})$ into $C^{*}({\mathcal{B}})$ (resp. from $C_{r}^{*}({\mathcal{A}})$ into $C_{r}^{*}({\mathcal{B}})$), cf. [27, Section 21.1]. We will later see (cf. Example 4.4) that in general, every morphism $(\varphi,T):{\mathcal{A}}\to{\mathcal{B}}$ gives rise to a *-homomorphism from $C^{*}({\mathcal{A}})$ into $C^{*}({\mathcal{B}})$; moreover, if we assume that ${\rm ker}(\varphi)$ is amenable, then we will also get a *-homomorphism from $C_{r}^{*}({\mathcal{A}})$ into $C_{r}^{*}({\mathcal{B}})$.
@@
Using [46, Lemma 4.1], it follows from this equality that if $T$ is ${\mathcal{B}}$-positive definite, then $L_{S}$ is positive in $\mathcal{L}_{B_{e}}(B_{\varphi(\mathbf{g})})$. As the map $R\mapsto L_{R}$ is a faithful *-homomorphism from $M_{\varphi(\mathbf{g})}({\mathcal{B}})$ into $\mathcal{L}_{B_{e}}(B_{\varphi(\mathbf{g})})$, this is equivalent to $S$ being positive in $M_{\varphi(\mathbf{g})}({\mathcal{B}})$. This shows the forward implication. The converse implication can be shown by reversing the arguments above.
@@
The first assertion follows readily by choosing all $g_{i}$’s to be equal to $e$ in the definition of positive definiteness. Next, pick an approximate unit $\{u_{i}\}_{i\in I}$ for $A$. Let $i\in I,g\in G,a\in A_{g}$. Choosing $g_{1}=e,g_{2}=g,a_{1}=u_{i}$ and $a_{2}=a$, and setting $\mathbf{g}=(e,g)$, we get that the matrix
@@
Moreover, since $T_{e}(u_{i}^{2})$ and $T_{e}(a^{*}a)$ are positive elements in $A$, it follows from [46, Lemma 5.2 (iii)] that
@@
Note that if ${\mathcal{A}}$ is unital (i.e., if $A_{e}$ is unital), replacing each $u_{i}$ with $1_{A_{e}}$, we don’t need to take any limit in the proof above, hence the boundedness assumption for each $T_{g}$ is not required in this case for the second and the third properties above to hold; the boundedness of each $T_{g}$ follows then readily from the third property, and could therefore be omitted in the definition of a positive definite bundle map.

### v2

@@ Abstract
We introduce a natural concept of positive definiteness for bundle maps between Fell bundles over (possibly different) discrete groups and describe several examples. Such maps induce completely positive maps between the associated full cross-sectional C*-algebras in a functorial way. Under the assumption that the kernel of the homomorphism connecting the groups under consideration is amenable, they also induce completely positive maps between the associated reduced cross-sectional C*-algebras. As an application, we define an approximation property for a Fell bundle over a discrete group which generalizes Exel’s approximation property and still implies the weak containment property. Both approximation properties coincide when the unit fibre is nuclear.
@@ 1 Introduction
The concept of positive definiteness for complex-valued functions on a group $G$ has a long history (it goes back to Toeplitz in 1911 when $G={\mathbb{Z}}$ and to Mathias in 1923 when $G={\mathbb{R}}$; for more details, see [60] and references therein). Its importance is ubiquitous in various areas of mathematics, in particular in connection with the development in the 1940’s and 50’s of the representation theory of locally compact groups and of the associated group $C^{*}$-algebras (see e.g. [22, 29, 30]). On the other hand, completely positive linear maps play a prominent rôle in the modern theory of operator algebras and in quantum information theory (cf. e.g. [57, 14, 34]). Haagerup discovered in 1978 that there is a very useful link between the two concepts. Namely, he proved in [32, Theorem 3.1] that given an action $\alpha$ of a locally compact group $G$ on a von Neumann algebra ${\cal M}$, every continuous positive definite function $\psi$ on $G$ gives rise to a normal completely positive linear map $M_{\psi}$ on the crossed product ${\cal M}\rtimes_{\alpha}G$ such that $M_{\psi}$ acts on each $g$-component by multiplication with $\psi(g)$. This link also holds in the context of group $C^{*}$-algebras and group von Neumann algebras, as noted by Haagerup in his seminal paper [33] and exploited in later works, such as [21] where continuous positive definite functions on $G$ are identified as multipliers of the Fourier algebra $A(G)$ of $G$. In the setting of discrete unital $C^{*}$-dynamical systems, a similar link was established in a series of paper [8, 9, 10], where broader concepts of multipliers were introduced. As a sample of related articles we mention [50, 51, 52, 35, 6, 40, 49].
@@
Given two $C^{*}$-algebras $A$ and $B$, it is often the case that one needs to construct completely positive linear maps from $A$ to $B$. When $A$ and $B$ belong to some class of $C^{*}$-algebras, one may hope to find some specific procedure achieving this. In this paper, we consider this problem for the class $\mathcal{F}$ (resp. $\mathcal{F}_{\rm r}$) of $C^{*}$-algebras which may be described as the full (resp. reduced) cross-sectional $C^{*}$-algebra $C^{*}({\mathcal{A}})$ (resp. $C_{r}^{*}({\mathcal{A}})$) of a Fell bundle ${\mathcal{A}}$ over a discrete group, as studied in [27, 28] (and e.g. in [46, 41]; see also [29, 30] for a more general treatment dealing with $C^{*}$-algebraic bundles, i.e., Fell bundles over locally compact groups). The class $\mathcal{F}$ (resp. $\mathcal{F}_{\rm r}$) is large as it contains all full (resp. reduced) twisted group $C^{*}$-algebras, and more generally all full (resp. reduced) $C^{*}$-crossed products associated to discrete twisted $C^{*}$-dynamical systems, even those for which the action of the group is only assumed to be partial [26]. Also, every $C^{*}$-algebra on which there exists a maximal (resp. normal) nondegenerate coaction by a discrete group belongs to $\mathcal{F}$ (resp. $\mathcal{F}_{\rm r}$), see [24, Proposition 4.2] (resp. [59, Corollary 3.9]). In particular, both classes include any $C^{*}$-algebra on which there exists a nondegenerate coaction by a discrete amenable group, e.g., any $C^{*}$-algebra on which there exists an action of a compact abelian group, such as Cuntz algebras and, more generally, (higher-rank) graph $C^{*}$-algebras. Note here that one may always choose to regard a $C^{*}$-algebra $B$ as the cross-sectional $C^{*}$-algebra of a Fell bundle over the trivial group, hence that our set up includes the case where no assumption is put on $B$.
@@
After a preliminary section (Section 2), our main findings are gathered in Section 3. To describe these, let us first assume that $A=C^{*}({\mathcal{A}})$ and $B=C^{*}({\mathcal{B}})$, where ${\mathcal{A}}$ is a Fell bundle over a discrete group $G$ and ${\mathcal{B}}$ is a Fell bundle over a discrete group $H$, and let $\varphi:G\to H$ be a homomorphism. Consider an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T:{\mathcal{A}}\to{\mathcal{B}}$, meaning that $T$ maps the fibre of ${\mathcal{A}}$ over any $g\in G$ into the fibre of ${\mathcal{B}}$ over $\varphi(g)$ in a linear and bounded way. Inspired by the definition of positive definiteness for a complex function on a group and the definition of complete positivity for a linear map between $C^{*}$-algebras, we introduce a natural notion of positive definiteness for such a bundle map, and show that $T$ is positive definite if and only if there exists a completely positive linear map $\Phi_{T}:A\to B$ which “extends” $T$ (see Theorem 3.11 for a more precise statement). Our proof relies on a reformulation of positive definiteness in terms of the matrix $C^{*}$ algebras associated to a Fell bundle by Abadie and Ferraro in [1], and builds on the Stinespring dilation theorem for Fell bundles due to Buss, Ferraro and Sehnem in [18, Appendix]. As we point out in Remark 3.19, the association ${\mathcal{A}}\mapsto C^{*}({\mathcal{A}}),\,(\varphi,T)\mapsto\Phi_{T}$ is functorial.
@@
Next, we consider the case where $A=C_{r}^{*}({\mathcal{A}})$ and $B=C_{r}^{*}({\mathcal{B}})$. It is not difficult to see that an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T$ has to be positive definite whenever there exists a completely positive linear map $M_{T}:A\to B$ which “extends” $T$. However, the converse statement can not hold in general. This may be deduced from a result in the recent book of Bekka and de la Harpe, cf. [11, Corollary 8.C.15], which says that the homomorphism $\varphi:G\to H$ extends to a *-homomorphism from $C_{r}^{*}(G)$ into $C_{r}^{*}(H)$ if and only if the kernel of $\varphi$ is an amenable subgroup of $G$. Using this result in combination with Fell’s absorption principle for Fell bundles and a generalization of another result of Buss, Ferraro and Sehnem ([18, Proposition 4.8]), we are able to show that the converse does hold if we also assume that $\ker(\varphi)$ is amenable, cf. Theorem 3.14. Again, the association ${\mathcal{A}}\mapsto C_{r}^{*}({\mathcal{A}}),(\varphi,T)\mapsto M_{T}$ is functorial in a suitable sense.
@@
In Section 4 we illustrate the usefulness of Theorem 3.11 and Theorem 3.14 in a series of examples, showing that these two results are not only generalizations of previous results of the same flavor, but also lead to new applications. In particular, we devote a section (Section 5) to discussing a new approximation property for a Fell bundle ${\mathcal{A}}$ over a discrete group. We call it the PD-approximation property since it requires the existence of a net of uniformly bounded, finitely supported positive definite bundle maps converging to the identity map on each fibre, in analogy with the well-known characterization of the amenability of a discrete group in terms of positive definite functions. The PD-approximation property for ${\mathcal{A}}$ is (formally) weaker than Exel’s approximation property [27, 28], but still implies the weak containment property (i.e., amenability of ${\mathcal{A}}$ in the sense of Exel), cf. Theorem 5.4. Moreover, when it is satisfied, then $C^{*}({\mathcal{A}})$ (resp. $C^{*}_{r}({\mathcal{A}})$) is nuclear if and only if the unit fibre of ${\mathcal{A}}$ is nuclear, cf. Theorem 5.6. Using current knowledge from [3], it follows that the PD-approximation property for ${\mathcal{A}}$ is equivalent to Exel’s approximation property when the unit fibre of ${\mathcal{A}}$ is assumed to be nuclear. However, the arguments used to deduce this fact give no clue whether it also holds when the unit fibre of ${\mathcal{A}}$ is not nuclear.
@@
When we were about to finish the first draft of the present article, an interesting preprint by Buss, Kwasniewski, McKee and Skalski appeared [19], where the existing theories of Fourier-Stieltjes algebras associated to groups [37], to groupoids [61, 54, 56] and to discrete twisted unital $C^{*}$-dynamical systems [10] is extended to the category of twisted actions by étale groupoids on $C^{*}$-bundles. Only a few of their results deal with the general case of Fell bundles over étale groupoids, hence with general Fell bundles over discrete groups, and there is little overlap between our papers.
@@ 2 Preliminaries
Definition 2.1.
([28, Definition 16.1]) A Fell bundle ${\mathcal{A}}=(A_{g})_{g\in G}$ over $G$ is a collection of Banach spaces (called fibers) satisfying the following properties (where ${\mathcal{A}}$ also denotes the disjoint union of the $A_{g}$’s). There is an associative multiplication map from ${\mathcal{A}}\times{\mathcal{A}}$ into ${\mathcal{A}}$ such that $A_{g}A_{h}\subseteq A_{gh}$, $(a,b)\mapsto ab$ is bilinear on $A_{g}\times A_{h}$ for all $g,h\in G$, and $\|ab\|\leq\|a\|\|b\|$ for all $a,b\in A$. Moreover, there is an involutive, anti-multiplicative map from ${\mathcal{A}}$ into itself such that $A_{g}^{*}=A_{g^{-1}}$ and $a\mapsto a^{*}$ is conjugate-linear, norm-preserving from $A_{g}$ into $A_{g^{-1}}$ for all $g\in G$. Finally, we have $\|a^{*}a\|=\|a\|^{2}$ and $a^{*}a\geq 0$ in the unit fibre $A_{e}$ (which is a $C^{*}$-algebra) for all $a\in{\mathcal{A}}$.
The vector space $C_{c}({\mathcal{A}})$ consisting of all finitely supported functions $f:G\to{\mathcal{A}}$ such that $f(g)\in A_{g}\text{ for all }g\in G$, becomes a *-algebra with respect to the operations defined by
@@
$\iota^{\mathcal{A}}(f)=\sum_{g\in G}\lambda^{\mathcal{A}}_{g}(f(g))\quad\text{ for all }f\in C_{c}({\mathcal{A}}),$
see [28, Proposition 17.9, (ii)].
@@
As shown in [28, Proposition 17.9, (iv)], the map $\widehat{j}^{{\mathcal{A}}}_{g}$ is isometric for every $g\in G$. Thus we get that
@@
$E_{g}^{\mathcal{A}}\big{(}\lambda^{\mathcal{A}}_{h}(a)\big{)}=\begin{cases}a,\quad\text{ if }g=h,\\ 0,\quad\text{ if}~g\neq h,\end{cases}$
cf. [28, Lemma 17.8]. For $x\in C_{r}^{*}({\mathcal{A}})$, $E_{g}^{\mathcal{A}}(x)$ may be thought of as the Fourier coefficient of $x$ at $g$. The map $E^{\mathcal{A}}:=E_{e}^{\mathcal{A}}$ is a faithful conditional expectation from $C_{r}^{*}({\mathcal{A}})$ onto $A_{e}$ (when one identifies $A_{e}$ with $\lambda_{e}^{\mathcal{A}}(A_{e})$), see [28, Propositions 17.13 and 19.3].
We will need the following lemma several times. It follows readily from [28, Lemma 17.2].
@@
is a *-algebra with respect to the natural operations, which can be equipped with a norm turning it into a $C^{*}$-algebra, cf. [1, Lemma 2.8]. We note that $M_{\mathbf{g}}({\mathcal{A}})$ can also be represented by adjointable operators on a right Hilbert $A_{e}$-module. Indeed, consider the right Hilbert $A_{e}$-module $A_{\mathbf{g}}:=A_{g_{1}^{-1}}\oplus\cdots\oplus A_{g_{n}^{-1}}$ obtained by taking the direct sum of the $A_{g_{i}}$’s (considered as right Hilbert $A_{e}$-modules), whose inner product is given by
@@ 3.1 Positive definite bundle maps
Considering the case where $G=H$ and $\varphi={\rm id}_{G}$, Exel defines a morphism from ${\mathcal{A}}$ to ${\mathcal{B}}$ to be an ${\mathcal{A}}$$-{\mathcal{B}}$ bundle map $T$ which is self-adjoint and multiplicative, and he proves that it gives rise to a *-homomorphism from $C^{*}({\mathcal{A}})$ into $C^{*}({\mathcal{B}})$ (resp. from $C_{r}^{*}({\mathcal{A}})$ into $C_{r}^{*}({\mathcal{B}})$), cf. [28, Section 21.1].
@@
Using [47, Lemma 4.1], it follows from this equality that if $T$ is ${\mathcal{B}}$-positive definite, then $L_{S}$ is positive in $\mathcal{L}_{B_{e}}(B_{\varphi(\mathbf{g})})$. As the map $R\mapsto L_{R}$ is a faithful *-homomorphism from $M_{\varphi(\mathbf{g})}({\mathcal{B}})$ into $\mathcal{L}_{B_{e}}(B_{\varphi(\mathbf{g})})$, this is equivalent to $S$ being positive in $M_{\varphi(\mathbf{g})}({\mathcal{B}})$. This shows the forward implication. The converse implication can be shown by reversing the arguments above.
Remark 3.5.
It will follow from Theorem 3.11 that $T$ is full (according to Definition 3.10). Similarly, if ${\rm ker}(\varphi)$ is amenable, Theorem 3.14 will imply that $T$ is also reduced. The associated maps $\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ and $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ are then actually *-homomorphisms, as is essentially well-known (cf. [28, Propositions 21.2 and 21.3], where the case $G=H$ and $\varphi={\rm id}_{G}$ is considered).
@@
The first assertion follows readily by choosing all $g_{i}$’s to be equal to $e$ in the definition of positive definiteness. Next, pick an approximate unit $\{u_{i}\}_{i\in I}$ for $A_{e}$. Let $i\in I,g\in G,a\in A_{g}$. Choosing $g_{1}=e,g_{2}=g,a_{1}=u_{i}$ and $a_{2}=a$, and setting $\mathbf{g}=(e,g)$, we get that the matrix
@@
Moreover, since $T_{e}(u_{i}^{2})$ and $T_{e}(a^{*}a)$ are positive elements in $A_{e}$, it follows from [47, Lemma 5.2 (iii)] that
@@
Note that if ${\mathcal{A}}$ is unital (i.e., if $A_{e}$ is unital), replacing each $u_{i}$ with $1_{A_{e}}$, we do not need to take any limit in the proof above, hence the boundedness assumption for each $T_{g}$ is not required in this case for the second and the third properties above to hold; the boundedness of each $T_{g}$ follows then readily from the third property (using that $T_{e}$ is completely positive, hence bounded), and could therefore be omitted in the definition of a positive definite bundle map.